Circular No. 6502

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1996bq, 1996br, AND 1996bs C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of two additional apparent supernovae on a panchromatic OCA film taken on Nov. 3.9 UT (limiting magnitude V ∼ 22) for confirming SN 1996bq (cf. IAUC 6500), which faded by ∼ 0.2 mag between Nov. 3.9 and 7.0.

SN 1996br is located at α = 2^{h}34^{m}53^{s}.53, δ = +41^{o}07′13′′.7 (equinox 2000.0), which is 1′′.6 east and 7′′.6 north of the center of UGC 2058; nothing was present at the position to V = 20.5 on OCA films taken on Oct. 12.1, to mag 22 on a 1995 OCA film, or on the original Palomar Sky Survey. Estimated magnitudes for SN 1996br: Nov. 3.9, V = 18.5; 7.0, 19.1. A nearby star star (V ∼ 18.0) has position end figures 52^{s}.03, 06′47′′.7.

SN 1996bs is located at α = 2^{h}44^{m}46^{s}.30, δ = +39^{o}29′26′′.5, which is 4′′.1 east and 0′′.3 north of the anonymous host galaxy. SN 1996bs superimposes a faint red condensation visible sometimes at mag V ∼ 20.5, but the 1995 OCA film (limiting mag ∼ 22) suggests that SN 1996bs is indeed a new star. Estimated magnitudes for SN 1996bs: Nov. 3.9, V ∼ 19.0; 7.0, 18.7. A nearby star (V ∼ 18.0) has position end figures 46^{s}.54, 34′′.7.

PSR B0656+14 A. Shearer, R. M. Redfern, G. Gorman, R. Butler, P. O’Kane, and A.

Golden, University College Galway; G. M. Beskin, S. I. Neizvestny, V. V. Neustroev, and V. L. Plokhotnichenko, Special Astrophysical Observatory, Nizhnij Arkhyz; and M. Cullum, European Southern Observatory, Garching, communicate: “We observed PSR B0656+14 on Jan. 14–15 with the 6-m telescope (+ TRIFFID camera and MAMA detector) and found optical pulsations in wavelength band B (mag B = 26.4). The significance of the result was 0.2 percent using the phase-independent Rayleigh statistics and 0.01 percent from studies of the Fourier power spectrum. The lightcurve showed a peak in phase with γ-ray data, and with a lag of 0.2 from the peak of the radio signal. Our pulsed magnitude is consistent with 100 percent of the signal being pulsed.”

SUPERNOVA 1996bk IN NGC 5308 Visual magnitude estimates (cf. IAUC 6491): Oct. 20.18 UT, 14.1 (J.

Carvajal, Teruel, Spain); Nov. 1.91, 14.0 (J. M. San Juan, Teruel, Spain); 2.81, 14.3 (Carvajal).

1996 November 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green